OpenStudy (anonymous):

The size of a computer monitor is given by the length of the diagonal of the screen. The ratio of the height to the width of wide-screen monitors is 9:16. Find expressions for the width and height of a wide-screen monitor in terms of the length of its diagonal. Possible answers: w=(16/sqrt 337) d , h=(9/sqrt 337) d w=(9/sqrt 337) d , h=(16/sqrt 337) d w=(16/sqrt 25) d , h=(9/sqrt 25) d w=(9/sqrt 25) d , h=(16/sqrt 25) d

OpenStudy (anonymous):

if w and h are 16 and 9 by pythagoras the diagonal^2: d^2 = 16^2 + 9^2 d = √337 so w = (16/√337)d and h = (9/√337)d

OpenStudy (anonymous):

Thank you!